OpenStudy (anonymous):

-7(3x-7)+21x>50 Always Sometimes Or Never ?

OpenStudy (anonymous):

@Brainybeauty

OpenStudy (anonymous):

\[-7(3x-7)+21x \ge 50\]

OpenStudy (anonymous):

When you solve the equation, -7(3x-7) +21x = -21x + 49 + 21x = 49 49 is less than 50 so the statement is false.

OpenStudy (anonymous):

@shawntaria